Lets assume that youre thinking about buying stock in West Coast Electronics. So far in your analysis, youve uncovered the

Round answers to two decimal places.
Answer #1

ANSWER

CALCULATION OF SECURITY EXPECTED RETURN USING IRR APPROACH

calculation of MPS after 3 years

MPS/EPS=11.6

MPS= 11.6*8.56

MPS= 99.296

formula to calculate IRR:

0 = P0 + P1/(1+IRR) + P2/(1+IRR)2 + P3/(1+IRR)3 + . . . +Pn/(1+IRR)n

where P0, P1, . . . Pn equals the cash flows in periods 1, 2, . . . n, respectively; and

IRR equals the project's internal rate of return

let us try IRR= 16%

0= -74.45 + 4.76 /(1+.16) + 4.76 /(1+.16)2 + (4.76+99.296)/(1+.16)3  

0 is not equal to 0.2

let us try IRR = 15.922%

here 0=0

hence IRR= 15.922%

DIVIDEND VALUATION MODEL

VALUE OF A SHARE OF STOCK=NEXT YEARS DIVIDEND

REQUIRED RATE OF RETURN- CONSTANT RATE OF GROWTH IN DIVIDENDS

=4.76

.15922-0

=29.896

CURRENTLY THE STOCK TRADES AT $74.45 WHEREAS ITS VALUE IS $29.896 . THE STOCK IS OVER VALUED AND HENCE NOT GOOD FOR INVESTMENT